OpenStudy (anonymous):

a large window air conditioner draws 11 amps of current. the unit runs for 8.0 hours during the course of a day. determine the quantity of charge that passes through the window AC during these 8.0 hours

OpenStudy (anonymous):

Current can be expressed as a function of Coulombs of charge in the following rate equation: \[I = {dQ \over dt}\] If we want to find the charge, Q, that passes over a given time, we need to integrate. \[dQ = I dt \rightarrow \int\limits dQ = \int\limits I dt \rightarrow Q = \int\limits_{t_f}^{t_i} I dt\]Solving the integrand assuming that current does not vary with time yields, \[Q = I \Delta t\]